OpenStudy (anonymous):

A drawer holds 4 red hats and 4 blue hats. What is the probability og getting exactly 3 red hats or exactly three blue hats when taking out 4 hats randomly out of the drawer and immediately returning every hat to the drawer before taking out the next? a.½ b. 1/8 c . ¼ d. 3/8

OpenStudy (anonymous):

there is conflict between a and b

OpenStudy (anonymous):

wait for a min...or till then can u write ur steps?

OpenStudy (anonymous):

finding 4 out of 8 (4/8*4/8*4/8*4/8)=1/16 1/16+1/16=2/16=1/8

OpenStudy (anonymous):

but why the conflict between a and b?

OpenStudy (anonymous):

when i google it i got another one i mean 1/2 thats why!!!!!!!!

OpenStudy (anonymous):

U get the answer on google also!

OpenStudy (anonymous):

yes on m4maths

OpenStudy (kropot72):

The binomial distribution applies to this problem. The probability of getting exactly three red hats or exactly three blue hats is given by: \[\left(\begin{matrix}4 \\ 3\end{matrix}\right)\times (0.5)^{3}\times 0.5=4\times (0.5)^{3}\times 0.5=?\]

OpenStudy (kropot72):

Required probability is 1/4.

OpenStudy (anonymous):

you are not going right

OpenStudy (kropot72):

@savi Do you understand the binomial distribution?

OpenStudy (anonymous):

i knew it but its a short time tricky question ,there are many way to solve it ,but i want an easy one

OpenStudy (anonymous):

4!/3! * (1/2)^4 = 1/4